OpenStudy (anonymous):

If y varies inversely as x, and y = 2 as x = 3, find y for the x-value of 1.

OpenStudy (anonymous):

y= Constant/x Constant = y times x Constant = 2 x 3 = 6 Therefore y=6/x For x-value of 1: y = 6/1 = 6.